# Leaf schemes and Hodge loci [^1] February 25, 2025

### v1

@@ 2.5 Pull-back Lamé equations
This section is based on many e-mail exchanges with S. Reiter who did most of the computation. I told him about a list of Lamé equations $g_{2}=0,g_{3}=1,B=0$ with at most one bad prime among all primes $\leq 200$ and a similar list for $g_{2}=1,g_{3}=0,B=0$ with at most two bad primes among all primes $\leq 100$. For example, the Lamé equation with $g_{2}=0,g_{3}=1,B=0,n=347/480$ , has only the bad prime $197$ among all primes $<=200$. The primes $p=2,3,5$ are not considered as they are factors of $480$.
@@
$(36z^{2}-9z)\frac{\partial^{2}}{\partial z^{2}}+(42z-6)\frac{\partial}{\partial z}-n(n+1),$
@@
$L:=(16z^{2}-4z)\frac{\partial^{2}}{\partial z^{2}}+(20z-3)\frac{\partial}{\partial z}-n(n+1)$
@@
S. Reiter also informed me of the classification of Pull-back Lamé equations with non Dihedral monodromy group by M. van Hoeij and R. Vidunas. Finally, the following question is rised by S. Reiter which is natural in view of computations done in [BvdW04]. For any Lamé equation with parameters $(n,B,g_{2},g_{3})$ and $k\in\mathbb{N}$ there is a $B_{k}$ such that Lamé equations with parameters $(n,B,g_{2},g_{3})$ and $(n+k,B_{k},g_{2},g_{3})$ are equivalent. For a suggestion on what an equivalence relation should mean see [BvdW04, page 3].
3 Local-global principle for leaf schemes
@@
One may ask whether imposing a certain Hodge class upon a generic member of an algebraic family of polarized algebraic varieties amounts to an algebraic condition upon the parameters, A. Weil in [Wei77].
@@ 3.1 Introduction
The present text arose from an attempt to combine Grothendieck-Katz $p$-curvature conjecture on linear differential equations with only algebraic solutions, and the Cattani-Deligne-Kaplan theorem on the algebraicity of Hodge loci. For our purpose we generalize Hodge loci into leaf schemes of foliations. This has first appeared in [Mov17b]. Foliations in the present text are given over finitely generated subrings of the field of complex numbers, and it makes sense to manipulate them modulo primes. We introduce conjectural criterions, 3.1 and 3.2, involving modulo $p$ manipulations of foliations which guarantee the algebraicity of leaf schemes.
@@
(27) ${{{\sum}}}_{\stackrel{{\scriptstyle a:I{}\to\mathbb{N}_{0},\ \beta:=\frac{1}{4}\left(\sum_{\alpha}a_{\alpha}\cdot\alpha+(1,1,1,1)\right),}}{{\beta_{i}\not\in\mathbb{Z},\ \beta_{0}+\beta_{1},\beta_{2}+\beta_{3}\in\mathbb{Z}}}}\frac{(-1)^{[\beta_{0}]+[\beta_{2}]}\langle\beta_{0}\rangle\langle\beta_{1}\rangle\langle\beta_{2}\rangle\langle\beta_{3}\rangle}{\prod_{\alpha\in I{}}a_{\alpha}!}\cdot\prod_{\alpha\in I{}}t_{\alpha}^{a_{\alpha}},$
where $I:=\{\alpha\in\mathbb{N}_{0}^{4}\ |\sum_{i=0}^{3}\alpha_{i}=4\}$, for a rational number $r$, $[r]$ is the integer part of $r$, that is $[r]\leq r<[r]+1$, $\{r\}=r-[r]$ and $\langle r\rangle=(r-1)(r-2)\cdots(\{r\})$. This function is a period of a holomorphic $2$-form on smooth surfaces of degree $4$ in $\mathbb{P}^{3}$ which is not algebraic. However, its zero locus is algebraic and parameterizes such surfaces with a line, for more details see Section 3.10.
@@
Throughout the text, ${\mathfrak{R}}$ is a finitely generated subring of $\mathbb{C}$ and ${\mathfrak{R}}_{\mathbb{Q}}:={\mathfrak{R}}\otimes_{\mathbb{Z}}\mathbb{Q}={\mathfrak{R}}[\frac{1}{2},\frac{1}{3},\frac{1}{5},\cdots]$ which is an infinitely generated subring of $\mathbb{C}$. We also take a finitely generated ring extension ${\mathfrak{R}}\subset{\mathsf{R}}$ and set ${\sf T}:={\rm Spec}({\mathsf{R}})$. By an ${\mathfrak{R}}$-scheme we simply mean the affine ${\mathfrak{R}}$-scheme ${\sf T}$. We consider an ${\mathfrak{R}}$-valued point of ${\sf T}$ which is nothing but an ${\mathfrak{R}}$-linear morphism ${t}:{\mathsf{R}}\to{\mathfrak{R}}$ and set $\mathfrak{m}_{{\sf T},{t}}:={\rm ker}({t})$. By definition, ${\cal O}_{\sf T}$ is just the ring ${\mathsf{R}}$, $\Omega^{1}_{\sf T}$ is the ${\cal O}_{\sf T}$-module of Kähler differentials (differential $1$-forms) in ${\sf T}$. For $f\in{\cal O}_{\sf T}$, $f({t}):={t}(f)$ is just the evaluation at ${t}$. The ${\cal O}_{\sf T}$-module $\Theta_{\sf T}$ of vector fields/derivations in ${\sf T}$ is the dual of the ${\cal O}_{\sf T}$-module $\Omega^{1}_{\sf T}$. We also look at a vector field ${\sf v}\in\Theta_{\sf T}$ as a derivation ${\sf v}:{\cal O}_{\sf T}\to{\cal O}_{\sf T},\ \ f\mapsto{\sf v}(f):={\sf v}(df)$. For an ${\mathfrak{R}}$-valued point ${t}$, we have also the well-defined ${\mathfrak{R}}$-linear map ${\sf v}(t):={t}\circ{\sf v}:\mathfrak{m}_{{\sf T},{t}}/\mathfrak{m}_{{\sf T},{t}}^{2}\to{\mathfrak{R}}$ which is called the evaluation of ${\sf v}$ at ${t}$. The dual ${{\bf T}}_{t}{\sf T}:=(\mathfrak{m}_{{\sf T},{t}}/\mathfrak{m}_{{\sf T},{t}}^{2})^{\vee}$ is the tangent space of ${\sf T}$ at ${t}$, and so, we have the evaluation map $\Theta_{\sf T}\to{{\bf T}}_{t}{\sf T}$ which is ${\mathfrak{R}}$-linear. As we do not need the language of sheaves, a sheaf on ${\sf T}$ is identified with the set of its global sections. We take a submodule $\Omega$ of $\Omega_{\sf T}$ and denote its dual by $\Theta:=\{{\sf v}\in\Theta_{\sf T}\ |\ {\sf v}(\Omega)=0\}$ which is a submodule of $\Theta_{\sf T}$.
We will assume that $\Omega$ is integrable in the strongest format, that is, $d\Omega\subset\Omega\wedge\Omega_{\sf T}$. Even though, we rarely use the integrability of $\Omega$, we will frequently use the notation ${\cal F}(\Omega)$ to denote the foliation induced by $\Omega$. From an algebraic point of view ${\cal F}(\Omega)$ is just $\Omega$ and nothing more. For any other subring $\check{\mathfrak{R}}$ of $\mathbb{C}$, we denote by ${\sf T}_{\check{\mathfrak{R}}}:={\sf T}\times_{\mathfrak{R}}{\rm Spec}(\check{\mathfrak{R}})$. The ring $\check{\mathfrak{R}}$ might be infinitely generated and the main example of this in this text is ${\mathfrak{R}}_{\mathbb{Q}}$. We denote by $({\sf T}^{\rm hol}_{\check{\mathfrak{R}}},{t})$ (resp. $({\sf T}^{\rm for}_{\check{\mathfrak{R}}},{t})$) the analytic (resp. formal) scheme underlying ${\sf T}$ and ${\cal O}_{{\sf T}^{\rm hol}_{\check{\mathfrak{R}}},{t}}$ (resp. ${\cal O}_{{\sf T}^{\rm for}_{\check{\mathfrak{R}}},{t}}$) is the ring of holomorphic functions in a neighborhood of $t$ (resp. formal power series) and with coefficients in $\check{\mathfrak{R}}$. We use the letter $L$ to denote a subscheme of $({\sf T}^{\rm hol},{t})$, that is, we have an ideal ${{\cal I}}\subset{\cal O}_{{\sf T}^{\rm hol},{t}}$ and ${\cal O}_{L}={\cal O}_{{\sf T}^{\rm hol},{t}}/{{\cal I}}$ (this might have zero divisors or nilpotent elements). Let $\Theta$ be a submodule of the ${\cal O}_{\sf T}$-module $\Theta_{\sf T}$. Its rank is the number $a\in\mathbb{N}$ such that $\wedge^{a+1}\Theta$ is a torsion sheaf and $\wedge^{a}\Theta$ is not. For ${\sf v}\in\Theta_{\sf T}$ we define the scheme ${\rm Sch}({\sf v}\in\Theta)$ given by the ideal generated by
$\left|\begin{matrix}{\sf v}(P_{1})&{\sf w}_{1}(P_{1})&{\sf w}_{2}(P_{1})&\cdots&{\sf w}_{a}(P_{1})\\ {\sf v}(P_{2})&{\sf w}_{1}(P_{2})&{\sf w}_{2}(P_{2})&\cdots&{\sf w}_{a}(P_{2})\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ {\sf v}(P_{a})&{\sf w}_{1}(P_{a})&{\sf w}_{2}(P_{a})&\cdots&{\sf w}_{a}(P_{a})\end{matrix}\right|,\ \ P_{1},P_{2},\cdots,P_{a}\in{\cal O}_{\sf T},\ \ {\sf w}_{1},{\sf w}_{2},\ldots,{\sf w}_{a}\in\Theta.$
@@
In geometric terms if ${\mathfrak{R}}$ is an algebraically closed field, ${\rm Sch}({\sf v}\in\Theta)$ is the loci of points ${t}$ in ${\sf T}$ such that the vector ${\sf v}({t})$ is in the ${\mathfrak{R}}$-vector space generated by ${\sf w}({t}),\ {\sf w}\in\Theta$. For two vector fields ${\sf v}$ and ${\sf w}$ the collinearness scheme ${\sf v}||{\sf w}$ is just ${\rm Sch}({\sf v}\in{\cal O}_{\sf T}{\sf w})$. For a primes $p$ which in not invertible in ${\mathfrak{R}}$, the ring ${\mathfrak{R}}_{p}:={\mathfrak{R}}/p{\mathfrak{R}}$ is non-zero, and we define ${\sf T}_{p}:={\sf T}\times_{\mathfrak{R}}{\rm Spec}({\mathfrak{R}}_{p})$ which is modulo $p$ reduction of ${\sf T}$. For a vector field ${\sf v}$ in ${\sf T}_{p}$, it is known that ${\sf v}^{p}$ is also a vector field, and this phenomena does not exist in ${\sf T}$.

### v2

@@ 2.5 Pull-back Lamé equations
This section is based on many e-mail exchanges with S. Reiter who did most of the computation. I told him about a list of Lamé equations $g_{2}=0,g_{3}=1,B=0$ with at most one bad prime among all primes $\leq 200$ and a similar list for $g_{2}=1,g_{3}=0,B=0$ with at most two bad primes among all primes $\leq 100$. For example, the Lamé equation with $g_{2}=0,g_{3}=1,B=0,n=347/480$ , has only the bad prime $197$ among all primes <=200<=200. The primes $p=2,3,5$ are not considered as they are factors of $480$.
@@
(2.8) $(36z^{2}-9z)\frac{\partial^{2}}{\partial z^{2}}+(42z-6)\frac{\partial}{\partial z}-n(n+1),$
@@
(2.9) $L:=(16z^{2}-4z)\frac{\partial^{2}}{\partial z^{2}}+(20z-3)\frac{\partial}{\partial z}-n(n+1)$
@@
By a linear transormation in $z$, one can see easily that the differential equation (2.8) and (2.9) can be transformed into the Gauss hypergeometric equation with $(a,b,c)=(\frac{-n}{6},\frac{n}{6}+\frac{1}{6},\frac{2}{3})$ and $(\frac{-n}{4},\frac{n}{4}+\frac{1}{4},\frac{3}{4})$, respectively. S. Reiter also informed me of the classification of Pull-back Lamé equations with non Dihedral monodromy group by M. van Hoeij and R. Vidunas. Finally, the following question is rised by S. Reiter which is natural in view of computations done in [Beukers2004]. For any Lamé equation with parameters $(n,B,g_{2},g_{3})$ and $k\in\mathbb{N}$ there is a $B_{k}$ such that Lamé equations with parameters $(n,B,g_{2},g_{3})$ and $(n+k,B_{k},g_{2},g_{3})$ are equivalent. For a suggestion on what an equivalence relation should mean see [Beukers2004, page 3].
We finish this section with some experiments with the Gauss hypergeometric equation:
(2.10) $z(1-z)y^{\prime\prime}+(c-(a+b+1)z)y^{\prime}-aby=0,\ \ a,b,c\in\mathbb{C}.$
For a random choice of parameters $a,b,c$ in the Gauss hypergeometric equation it seems that the $p$-curvature is always non-zero except for a finite number of primes $p$. For example, for $(a,b,c)=(\frac{1}{2},\frac{1}{2},1)$ and $p\not=2$, the $p$-curvature seems to never vanish. We have checked this for primes $\leq 150$.
LIB "foliation.lib";
ring r=(0,z),x,dp; matrix lde[1][3]; int i; int j;
number a=1/2; number b=1/2; number c=1;
int ub=150; number parm=z;
          lde=-a*b, c-(a+b+1)*z, z*(1-z);
          BadPrD(lde, parm, ub);
From another side in 2.4 we have got the Gauss hypergeometric equations with $(a,b,c)=(\frac{-n}{6},\frac{n}{6}+\frac{1}{6},\frac{2}{3})$ and $(\frac{-n}{4},\frac{n}{4}+\frac{1}{4},\frac{3}{4})$, for which we can experimentally observe that the number of both vanishing and non-vanishing curvature is infinite.
LIB "foliation.lib";
ring r=(0,z),x,dp;matrix lde[1][3]; int i; int j;
number a; number b; number c=2/3;
int ub=50; number parm=z; number N=100;
   for (i=1; i<=N; i=i+1)
       for (j=1; j<=N; j=j+1)
          a=-i/number(6*j);  b=-i/number(6*j)+1/6;
         lde=-a*b, c-(a+b+1)*z, z*(1-z);
          BadPrD(lde, parm, ub);
2.6 Differential equations and density of primes
The discussion in this section motivates us to define the density of primes for which the $p$-curvature vanishes. Recall that we have two notions of density. For a subset $S$ of primes, we say that $S$ has the natural, respectively analytic or Dirichlet, density $\delta$ if the following limit exists and it equals to $\delta$:
$\lim_{x\to\infty}\frac{\#\{p\in S\ |\ p\leq x\}}{\#\{p\hbox{ prime}\ |\ p\leq x\}}\ \ \hbox{ respectively }\ \ \lim_{s\to 1}\frac{\sum_{p\in S}\frac{1}{p^{s}}}{\sum_{p\hbox{ prime}}\frac{1}{p^{s}}}.$
If a set of primes has a natural density, then it has also an analytic density and the two densities are equal, but the converse is false, see [LenstraStevenhagen]. For a linear differential equation $\frac{\partial y}{\partial z}={\sf A}(z)y$, we define its curvature denisty (natural or analytic), the density of primes $p$ such that its $p$-curvature vanishes. The density of a finite number of primes (resp. all primes except for a finite number) is zero (resp. one). The converse is not necessarily true, see the mathoverflow link. One might try to formulate a stronger version of the Grothendieck-Katz conjecture speculating that if the curvature density of a differential equation is one then all its solutions are algebraic. The only well-known result about curvature density is the following:
Proposition 2.5.
Let $a$ be an algebraic number. The natural and analytic curvature density of $y^{\prime}=\frac{a}{z}y,\ a\in\mathbb{C}^{*}$ is
$\frac{1}{\#{\rm Gal}({\sf k}/\mathbb{Q})},$
where $\sf k$} is the splitting field of the minimal polynomial of $a$.
Proof.
Let $f(x)$ be the minimal polynomial of $a$. If the $p$-curvature of $y^{\prime}=\frac{a}{z}y$ is zero then $a^{p}-a=0$ in $\mathbb{Z}[a]/p\mathbb{Z}[a]$. This implies that $f$ divides $x^{p}-x$ in $\mathbb{F}_{p}[x]$. In particular, $f$ modulo $p$ is a product of linear equations. The proposition is now a direct consequence of the Frobenius density theorem: Let $f\in\mathbb{Q}[x]$ be a polynomial. The density of primes $p$ such that
$f\ {\rm mod}\ p=f_{1}f_{2}\cdots f_{t},\ f_{i}\in\mathbb{F}_{p}[x]\hbox{ irreducible },\ \deg(f_{i})=n_{i},\$
is equal to
(2.11) $\frac{\#\{\sigma\in{\rm Gal}({\sf k}/\mathbb{Q})\ |\sigma\hbox{ with cycle pattern }n_{1},n_{2},\ldots,n_{t}\ \}}{\#{\rm Gal}({\sf k}/\mathbb{Q})}.$
see [LenstraStevenhagen]. For the application of this theorem we consider the case $n_{1}=n_{2}=\cdots=n_{t}=1$
Frobenius’ density theorem is a particular case of Chebotarev’s density theorem. The author tried to put these theorems in terms of curvature density of differential equations but there was no succces beyond 2.5. Even for the Gauss hypergeometric function obtained in 2.4, the author is not aware of any result regarding the density of vanishing $p$-curvatures.
Each computation of the present chapter is done in less than an hour, and it turns out that if we let our codes run for a day or more we can get more accurate density statements, for instance, for the Lamé equation with $(n,B,g_{2},g_{3})=(\frac{5}{87},0,0,1)$ after a day of computation, the author was able to compute all good ($p$-curvature zero) and bad primes $\leq 797$. After the first $50$ primes the first digit of the density of good primes seems to stabilize in $0.7$, however, the second digit does not seem to stabilize. The density of good primes $p\leq 797$ is $0.71428$. Similar statements seem to be true for the Lamé equation with $(n,B,g_{2},g_{3})=(\frac{12}{89},0,0,1)$. For the computer codes and data see the author’s webpage.
Chapter 3 Local-global principle for leaf schemes
@@ 2.6 Differential equations and density of primes
One may ask whether imposing a certain Hodge class upon a generic member of an algebraic family of polarized algebraic varieties amounts to an algebraic condition upon the parameters, A. Weil in [Weil1977].
@@ 3.1 Introduction
The present text arose from an attempt to combine Grothendieck-Katz $p$-curvature conjecture on linear differential equations with only algebraic solutions, and the Cattani-Deligne-Kaplan theorem on the algebraicity of Hodge loci. For our purpose we generalize Hodge loci into leaf schemes of foliations. This has first appeared in [GMCD-NL]. Foliations in the present text are given over finitely generated subrings of the field of complex numbers, and it makes sense to manipulate them modulo primes. We introduce conjectural criterions, 3.1 and 3.2, involving modulo $p$ manipulations of foliations which guarantee the algebraicity of leaf schemes.
@@
(3.1) ${{{\sum}}}_{\stackrel{{\scriptstyle a:I{}\to\mathbb{N}_{0},\ \beta:=\frac{1}{4}\left(\sum_{\alpha}a_{\alpha}\cdot\alpha+(1,1,1,1)\right),}}{{\beta_{i}\not\in\mathbb{Z},\ \beta_{0}+\beta_{1},\beta_{2}+\beta_{3}\in\mathbb{Z}}}}\frac{(-1)^{[\beta_{0}]+[\beta_{2}]}\langle\beta_{0}\rangle\langle\beta_{1}\rangle\langle\beta_{2}\rangle\langle\beta_{3}\rangle}{\prod_{\alpha\in I{}}a_{\alpha}!}\cdot\prod_{\alpha\in I{}}t_{\alpha}^{a_{\alpha}},$
where $I:=\{\alpha\in\mathbb{N}_{0}^{4}\ |\sum_{i=0}^{3}\alpha_{i}=4\}$, for a rational number $r$, $[r]$ is the integer part of $r$, that is $\leq r<[r]+1, {r}=r$$\{r\}$=r-[r] and $\langle r\rangle=(r-1)(r-2)\cdots(\{r\})$. This function is a period of a holomorphic $2$-form on smooth surfaces of degree $4$ in $\mathbb{P}^{3}$ which is not algebraic. However, its zero locus is algebraic and parameterizes such surfaces with a line, for more details see Section 3.10.
@@
Throughout the text, $\mathfrak{R}$} is a finitely generated subring of $\mathbb{C}$ and ${\mathfrak{R}}_{\mathbb{Q}}:={\mathfrak{R}}\otimes_{\mathbb{Z}}\mathbb{Q}={\mathfrak{R}}[\frac{1}{2},\frac{1}{3},\frac{1}{5},\cdots]$ which is an infinitely generated subring of $\mathbb{C}$. We also take a finitely generated ring extension ${\mathfrak{R}}\subset{\mathsf{R}}$ and set ${\sf T}:={\rm Spec}({\mathsf{R}})$. By an $\mathfrak{R}$}-scheme we simply mean the affine $\mathfrak{R}$}-scheme $\sf T$}. We consider an $\mathfrak{R}$}-valued point of $\sf T$} which is nothing but an $\mathfrak{R}$}-linear morphism ${t}:{\mathsf{R}}\to{\mathfrak{R}}$ and set $\mathfrak{m}_{{\sf T},{t}}:={\rm ker}({t})$. By definition, ${\mathcal{O}}_{\sf T}$ is just the ring $\mathsf{R}$}, $\Omega^{1}_{\sf T}$ is the ${\mathcal{O}}_{\sf T}$-module of Kähler differentials (differential $1$-forms) in $\sf T$}. For $f\in{\mathcal{O}}_{\sf T}$, f(t):=t(f)f({t}):={t}(f) is just the evaluation at t{t}. The ${\mathcal{O}}_{\sf T}$-module $\Theta_{\sf T}$ of vector fields/derivations in $\sf T$} is the dual of the ${\mathcal{O}}_{\sf T}$-module $\Omega^{1}_{\sf T}$. We also look at a vector field ${\sf v}\in\Theta_{\sf T}$ as a derivation ${\sf v}:{\mathcal{O}}_{\sf T}\to{\mathcal{O}}_{\sf T},\ \ f\mapsto{\sf v}(f):={\sf v}(df)$. For an $\mathfrak{R}$}-valued point t{t}, we have also the well-defined $\mathfrak{R}$}-linear map ${\sf v}(t):={t}\circ{\sf v}:\mathfrak{m}_{{\sf T},{t}}/\mathfrak{m}_{{\sf T},{t}}^{2}\to{\mathfrak{R}}$ which is called the evaluation of $\sf v$} at t{t}. The dual ${{\bf T}}_{t}{\sf T}:=(\mathfrak{m}_{{\sf T},{t}}/\mathfrak{m}_{{\sf T},{t}}^{2})^{\vee}$ is the tangent space of $\sf T$} at t{t}, and so, we have the evaluation map $\Theta_{\sf T}\to{{\bf T}}_{t}{\sf T}$ which is $\mathfrak{R}$}-linear. As we do not need the language of sheaves, a sheaf on $\sf T$} is identified with the set of its global sections. We take a submodule $\Omega$ of $\Omega_{\sf T}$ and denote its dual by $\Theta:=\{{\sf v}\in\Theta_{\sf T}\ |\ {\sf v}(\Omega)=0\}$ which is a submodule of $\Theta_{\sf T}$.
[…]
We use the letter $L$ to denote a subscheme of $({\sf T}^{\rm hol},{t})$, that is, we have an ideal ${{\mathcal{I}}}\subset{\mathcal{O}}_{{\sf T}^{\rm hol},{t}}$ and ${\mathcal{O}}_{L}={\mathcal{O}}_{{\sf T}^{\rm hol},{t}}/{{\mathcal{I}}}$ (this might have zero divisors or nilpotent elements). Let $\Theta$ be a submodule of the ${\mathcal{O}}_{\sf T}$-module $\Theta_{\sf T}$. Its rank is the number $a\in\mathbb{N}$ such that $\wedge^{a+1}\Theta$ is a torsion sheaf and $\wedge^{a}\Theta$ is not. For ${\sf v}\in\Theta_{\sf T}$ we define the scheme ${\rm Sch}({\sf v}\in\Theta)$ given by the ideal generated by
$\left|\begin{matrix}{\sf v}(P_{1})&{\sf w}_{1}(P_{1})&{\sf w}_{2}(P_{1})&\cdots&{\sf w}_{a}(P_{1})\\ {\sf v}(P_{2})&{\sf w}_{1}(P_{2})&{\sf w}_{2}(P_{2})&\cdots&{\sf w}_{a}(P_{2})\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ {\sf v}(P_{a})&{\sf w}_{1}(P_{a})&{\sf w}_{2}(P_{a})&\cdots&{\sf w}_{a}(P_{a})\end{matrix}\right|,\ \$
$P_{1},P_{2},\cdots,P_{a}\in{\mathcal{O}}_{\sf T},\ \ {\sf w}_{1},{\sf w}_{2},\ldots,{\sf w}_{a}\in\Theta.$
@@
In geometric terms if $\mathfrak{R}$} is an algebraically closed field, ${\rm Sch}({\sf v}\in\Theta)$ is the loci of points t{t} in $\sf T$} such that the vector ${\sf v}({t})$ is in the $\mathfrak{R}$}-vector space generated by ${\sf w}({t}),\ {\sf w}\in\Theta$. For two vector fields $\sf v$} and $\sf w$} the collinearness scheme ${\sf v}||{\sf w}$ is just ${\rm Sch}({\sf v}\in{\mathcal{O}}_{\sf T}{\sf w})$. For a primes $p$ which in not invertible in $\mathfrak{R}$}, the ring ${\mathfrak{R}}_{p}:={\mathfrak{R}}/p{\mathfrak{R}}$ is non-zero, and we define ${\sf T}_{p}:={\sf T}\times_{\mathfrak{R}}{\rm Spec}({\mathfrak{R}}_{p})$ which is modulo $p$ reduction of $\sf T$}. For a vector field $\sf v$} in ${\sf T}_{p}$, it is known that ${\sf v}^{p}$ is also a vector field, and this phenomena does not exist in $\sf T$}.